3. An aerospace company has submitted bids on two separate government defense contracts, A and B. The company feel it has a 50% chance of winning contract A and a 40% chance of winning contract B. Furthermore, it believes that winning contract A is independent of winning contract B.
a. What is the probability that the company will win both contracts?
b. What is the probability that the company will win either of these contracts?

Answer by ikleyn(52781)

.
An aerospace company has submitted bids on two separate government defense contracts, A and B.
The company feel it has a 50% chance of winning contract A and a 40% chance of winning contract B.
Furthermore, it believes that winning contract A is independent of winning contract B.
a. What is the probability that the company will win both contracts?
b. What is the probability that the company will win either of these contracts?
~~~~~~~~~~~~~~~

(1)  P(winning both contracts) = P(A)*P(B) = 0.5*0.4 = 0.2.    ANSWER to (a)


(2)  P(winning at least one contract) = P(A) + P(B) - P(winning both) = 0.5 + 0.4 - 0.2 = 0.7.    ANSWER to (b)

Solved.